# From Evaluation to Enhancement: Large Language Models for Zero-Knowledge Proof Code Generation

### v1

@@ 4. Augmenting LLMs for ZK Code Generation
Our study reveals a gap between LLMs’ surface-level knowledge of ZK DSLs and their ability to construct correct gadgets or complete programs. Models can recall syntax and library references but struggle to translate this knowledge into valid constraint systems. To bridge this gap, we introduce ZK-Coder, an agentic framework for end-to-end ZK code generation. We first summarize the design principles derived from our findings, then present the system.
@@ 4.1. Design Principles Learned from the Findings
Findings ① and ② show that LLMs possess a strong knowledge of ZK DSL syntax, semantics, and standard libraries, achieving near expert-level accuracy even on advanced features. This reflects their absorption of publicly available documentation and indicates their potential for ZK code generation. Despite this, limitations remain that motivate our further exploration.
In Findings ③ and ④, we observe that this knowledge does not translate into reliable gadget-level generation. Generated code frequently suffers from typing errors, signal mismanagement, and faulty constraint enforcement, reflecting the scarcity of training examples with concrete gadget implementations. To bridge this gap between knowing the language and encoding correct programs, we introduce a sketch language in ZK-Coder that abstracts ZKP’s underlying constraint system. To further link them to curated hints, ZK-Coder leverages retrieval-augmented generation (RAG) (Gao et al., 2023) to guide gadget usage. Loosely speaking, RAG trades brittle memorization for grounding: once the needed gadgets are identified, retrieval supplies exact usage patterns that collapse the search space and curb gadget hallucinations. This strategy explicitly mirrors how developers consult a codebook and enables composition of verified snippets for rare gadgets that pretraining scarcely covers.
Finally, Findings ③ and ⑤ highlight end-to-end program generation as the most challenging setting, where generated programs often fail due to syntactic errors or semantic mismatches. These failures directly undermine the reliability and necessitate mechanisms to improve validity and faithfulness. To mitigate these failures, ZK-Coder introduces a correctness-centric interactive loop. The LLM proposes a program, queries a compiler as an oracle to check syntactic validity, and, once compilable, executes test cases as an oracle to check intended semantics. By observing diagnostics and counterexamples, ZK-Coder repairs its code and iterates the “generate-compile-test-repair” loop until both syntactic and semantic checks pass. This interaction transforms failures into actionable feedback and guides candidates toward valid, faithful programs.
@@ 4.2. ZK-Coder
Based on these design principles, we developed ZK-Coder, illustrated in Fig. 10. ZK-Coder augments LLMs with a structured pipeline that bridges natural-language problem descriptions and executable ZK programs. It consists of three main stages: ➀ constraint formulation, ➁ constraint-guided analysis and retrieval, and ➂ interactive generation and repair. In ➀, problem statements are translated into lightweight sketch language. Then, in ➁, the sketch is analyzed to extract its constraints, which are used to retrieve relevant usage patterns on gadgets from a curated knowledge base. Finally, in ➂, the LLM generates the target ZK code using the retrieved hints, interactively repairing it with the help of syntax and semantic oracles until correctness is achieved. We detail the design of these stages in this section, and present an illustrative workflow with concrete code examples in the Supplementary Material.
@@
The first stage of ZK-Coder reasons and translates a natural-language problem description into a structured representation of ZKP programs. In accordance to our observations in Sec. 3.4 as well as prior work on sketching and planning for code generation (Wang et al., 2024; Deng et al., 2024; Li et al., 2023; Jiang et al., 2024a; Wen et al., 2024a), we introduce a lightweight intermediate language, the ZK Sketch Language (ZKSL; Fig. 11) that shares a similar essence with Python and Z3 ecosystem while abstracting away ZK-specific details. In ZKSL, a Program consists of a single verify function with typed parameters and a sequence of statements (Stmt); statements include assignments, assertions via engine.add($\Phi$), bounded loops and conditionals, and an optional return. Formulas $\Phi$ are built from relational atoms $E\bowtie E$ and gadget calls $G(\cdot)$. Its primitives (literals, arithmetic operators, etc.) align with the gadgets used in Sec. 3.1.1.
ZK-Coder prompts LLMs to generate ZKSL sketches from problem specifications. Then, to ensure well-formedness, ZK-Coder incorporates a checker that validates sketches on syntactic validity, type/arity consistency, and closure. If a sketch is invalid, feedback is provided to the LLM to refine its output. Since ZKSL is ZK-agnostic, it isolates constraint reasoning from ZK-specific syntax and semantics and reduces the generation complexity.
4.2.2. Constraint-Guided Analysis and Retrieval.
In addition to sketching and planning with ZKSL, RAG has been shown powerful to aid LLM-based code generation in low-resource, domain-specific languages (Joel et al., 2024). After obtaining a ZK-agnostic sketch, ZK-Coder analyzes its constraints and retrieves ZK-specific implementation guidance, bridging the abstract ZKSL specification and the concrete gadget implementations required in particular ZK DSLs.
We traverse the ZKSL sketch, collecting each constraint introduced via engine.add($\cdot$). Operators are canonicalized: aliases are normalized (e.g., >= is normalized to GreaterEqThan) and operator variants are distinguished by arity and types (e.g., XOr: Bool $\times$ Bool $\rightarrow$ Bool vs. XOr: Field $\times$ Field $\rightarrow$ Field). Non-constraint constructs (e.g., folded constants, loop statements or conditional guards that emit no constraints) are discarded. The result is an ordered, typed set of constraint primitives corresponding to the required gadgets.
To build a knowledge base for retrieval, we curate a collection of gadget and their usage examples either from the standard libraries of ZK DSLs or from verified implementations. With the knowledge base, ZK-Coder performs constraint-guided retrieval to fetch relevant usage patterns on gadgets for each constraint primitive. Retrieval enforces exact operator/type/arity matching, and hints are emitted in the ZKSL order to preserve program flow during generation. In this way, ZK-Coder effectively grounds generation in precise constraint analysis and structured retrieval.
4.2.3. Interactive Generation and Repair.
With the sketch generated in ➀ and the hints retrieved in ➁, ZK-Coder proceeds to the final stage of end-to-end code generation. As pointed out in prior work (Joel et al., 2024; Wang et al., 2022b; Bi et al., 2024; Skreta et al., 2023), introducing a feedback loop that leverages compilers and test executors as oracles can substantially improve code generation reliability. Motivated by this, ZK-Coder implements an interactive generate-repair loop to ensure both syntactic validity and semantic correctness of the generated ZK programs. The loop alternates between code generation, compilation checks, and execution-based testing, iteratively refining the program until it meets all correctness criteria.
Given the ZKSL sketch specification and constraint-specific hints, the LLM produces a candidate program. Then, ZK-Coder submits the program to the ZKP compiler (i.e., Circom or Noir) for compilation. If ZK-Coder observes compilation errors (e.g., syntax errors, type mismatches, or illegal signal operations), it extracts the diagnostics and feeds them back to the LLM for repair. The LLM refines the program based on this feedback, producing a new candidate that is again submitted for compilation. This process repeats until a compilable program is obtained. With a compilable program, ZK-Coder generates a set of test cases and validates the program on these cases on the fly. If any test case fails, ZK-Coder collects the counterexample and feeds it back to the LLM for further repair. This loop continues until the program successfully passes all test cases or the retry limit is reached. This interactive approach in ZK-Coder converts failures into constructive feedback and augments the LLM’s ability to produce reliable ZK programs.
@@ 5. Evaluation
We evaluate ZK-Coder on the task of generating zero-knowledge programs from natural language specifications. Our evaluation is structured around the following research questions, with a focus on end-to-end generation. To clarify, we do not re-evaluate language knowledge (RQ1) or gadget-level coding (RQ2): the former is established in Sec. 3.3, and the latter is handled by our retrieval mechanism in Sec. 4.2.2. Given that said, we include an ablation study RQ5.
@@
  • RQ6 (Robust Generalization): Beyond the standard benchmark, adapted HumanEval, how well does ZK-Coder transfer to newer and more challenging tasks?
  • RQ7 (Failure Analysis): When ZK-Coder fails, what are the common failure patterns, and how might they inform future research?
@@ 5.1. Experimental Setup
Our experiments are organized according to the research questions outlined above.
For RQ4 (End-to-End Performance), we evaluate on the adapted HumanEval benchmark from ZK-Eval, using the two strongest OpenAI’s reasoning models from our measurement study: GPT-o4-mini and GPT-o3. This comparison establishes how well ZK-Coder outperforms baseline methods in generating correct and provable ZK programs.
For RQ5 (Ablation Study), we conduct an ablation study on GPT-o4-mini, selectively disabling each design component of ZK-Coder, including the sketch language, constraint-guided retrieval, and the interactive repair loop. This aims to quantify their individual impact on overall performance.
@@
All experiments are conducted under a fixed repair budget of up to 8 iterations for syntax checking and repair, and 1 iteration for semantic checking and comparison. Each task is executed 10 times to compute Pass@1 accuracy, ensuring statistical stability. The evaluation is conducted on a Linux (Ubuntu 22.04 LTS) server equipped with 256 GiB of RAM.
@@ 5.2. RQ4: End-to-End Performance
We evaluate the end-to-end performance of ZK-Coder. To the best of our knowledge, since no prior work directly addresses zero-knowledge program generation from natural language, we establish a simple but meaningful baseline: directly prompting the model to produce circuit code from the task description and input specification, without any sketching, retrieval, or repair. This setting captures the raw capability of LLMs and provides a lower bound against which the contributions of ZK-Coder can be measured.
During the evaluation, we meticulously document the performance for the pipeline in three key stages: (1) Sketch correctness measures whether the initial ZK sketch produced from the natural language description is correct by passing all hidden test cases; (2) Repair pass rate is conditioned on all correctly generated sketches, and reports the fraction of these that can be successfully repaired by the interactive repair loop; (3) Program correctness is conditioned on all successfully repaired programs, and measures the proportion that pass all hidden test cases. The final overall accuracy aggregates across the full pipeline, reflecting end-to-end program generation accuracy. In addition, we report the average token cost for each model, which quantifies the spending.
@@
Circom GPT-o4-mini 94.12% 91.72% 96.59% 83.38% 2350.31 17.35% 436.98
GPT-o3 97.30% 94.73% 96.82% 89.23% 2253.29 20.29% 433.35
Noir GPT-o4-mini 94.12% 97.97% 97.66% 90.05% 648.95 32.21% 152.18
GPT-o3 97.21% 97.43% 99.38% 94.12% 765.46 27.94% 195.34
Table 3 summarizes the comparative performance of ZK-Coder and the baseline across both Circom and Noir. Across all settings, ZK-Coder achieves substantial gains over the baseline. On Circom, GPT-o4-mini and GPT-o3 reach overall pass rates of 83.38% and 89.23%, respectively, compared to only 17.35% and 20.29% for the baseline. Noir shows a similar trend, with GPT-o4-mini reaching 90.05% and GPT-o3 achieving the highest overall rate of 94.12%, far outperforming the corresponding baseline results of 32.21% and 27.94%.
The per-stage breakdown highlights the robustness of the pipeline. Sketch correctness exceeds 94% across all configurations, showing that initial translation into ZK sketches is highly reliable. The repair loop maintains strong success rates, above 91% in Circom and above 97% in Noir, ensuring that most sketches can be generated into valid programs. Final program correctness surpasses 96% in all cases, confirming that the integration of sketching, repair, and semantic validation yields consistently executable and provable programs.
We also note that ZK-Coder incurs a higher token cost compared to the baseline, which is expected since our agentic pipeline includes sketching, RAG and interactive repair. However, the cost remains modest in practice, with each task requiring fewer than 3,000 tokens on average, corresponding to less than 0.1 USD per task, making the gains in correctness highly cost-effective.
@@ 5.3. RQ5: Ablation Study
Figure 12. Accuracy Comparison of ZK-Coder Ablation Variants.
@@
  • No RAG (V1): Disables constraint-guided retrieval augmented generation, forcing the model to rely solely on its internal knowledge.
  • No Sketch (V2): Removes the intermediate ZK sketch reasoning and directly prompts the model to output full programs with RAG and the interactive repair loop.
  • No Syntax Repair (V3): Eliminates the interactive repair loop for syntax errors.
  • No Semantic Repair (V4): Retains syntax repair but disables semantic repair.
  • Similarity-based RAG (V5): Replaces our constraint-guided retrieval with a naive similarity-based retrieval using surface-level embedding similarity.
  • Only Repair (V6): Omits both retrieval and sketching, relying exclusively on the repair loop applied to direct generations.
Figure 12 presents the pass@1 accuracy for each ablation variant across Circom and Noir. The results show that every component plays a critical role in the overall success of ZK-Coder. Removing RAG (V1) reduces accuracy from 83.38% to 67.21% on Circom, and from 90.05% to 89.22% on Noir. Disabling the sketch reasoning stage (V2) causes further degradation (71.32% on Circom, 87.50% on Noir), highlighting the value of sketching as a scaffold for task understanding and constraint reasoning. The most severe drop arises from removing syntax repair (V3), with accuracy falling to 35.00% on Circom and 41.39% on Noir, demonstrating that interactive repair is indispensable. Disabling semantic repair (V4) also leads to measurable reductions (81.18% on Circom, 87.99% on Noir), confirming that semantic validation contributes to robustness. Substituting constraint-guided retrieval with a naive similarity-based strategy (V5) yields weaker performance (69.87% on Circom, 89.41% on Noir), suggesting that constraint-guided retrieval is more effective than surface-level matching. Finally, the “Only Repair” variant (V6) performs poorly, especially on Circom (59.12%), indicating that repair alone cannot compensate for the absence of sketching and retrieval.
Answer to RQ5: The ablation study shows that the strong performance of ZK-Coder relies on the synergy of sketching, constraint-guided RAG, and interactive repair. Removing or weakening any of these components causes substantial accuracy drops.

### v2

@@ 4. Augmenting LLMs for ZK Code Generation
Based on the findings and gaps revealed from our empirical study in § 3, we introduce ZK-Coder, an agentic framework to augment end-to-end ZK code generation. We first summarize the design principles derived from our findings, then present the system.
@@ 4.1. Design Principles Learned from the Findings
Findings ① and ② show that LLMs possess strong knowledge of ZK languages’ syntax, semantics, and standard library APIs, achieving near expert-level accuracy. Overall, we interpret that knowledge has been internalized during contemporary LLM’s pretraining stage, eliminating the need for grammar prompting or documentation access in our augmentation. However, important challenges remain that motivate our further exploration.
In Findings ③ and ④, we observe that this knowledge does not translate into reliable primitive-level generation. Generated code frequently suffers from typing errors, signal mismanagement, and faulty constraint enforcement. Finding ⑤ also reveals these failures in end-to-end generation. To bridge this gap, ZK-Coder introduces a sketch layer that decouples abstract constraint modeling from low-level DSL implementation. By first formalizing relations in this intermediate representation, the framework can precisely map identified primitives to verified circuit templates and usage hints via retrieval-augmented generation (RAG) (Gao et al., 2023). Loosely speaking, this trades brittle memorization for grounding. Once the logical “what” is defined in the sketch, retrieval provides the exact “how” for DSL-specific wiring, collapsing the search space and enabling the reliable composition of even rare algebraic snippets.
Finally, Findings ③ and ⑤ highlight end-to-end program generation as the most challenging setting, where generated programs often fail due to constraint-misspecification or under/over-constrained logic. These failures directly undermine the reliability and necessitate mechanisms to improve validity and faithfulness. To mitigate these failures, ZK-Coder introduces an interactive refinement loop. The LLM proposes a program, queries the compiler as an oracle to check constraint specification validity, and, once passes, self-supplies acceptance/rejection test cases to validate relation faithfulness. By observing diagnostics and counterexamples, ZK-Coder repairs its code and iterates the “generate-compile-test-repair” loop until these checks pass. This interaction transforms failures into actionable feedback and guides candidates toward valid, faithful programs.
@@ 4.2. ZK-Coder
Based on these design principles, we developed ZK-Coder, illustrated in Fig. 7. ZK-Coder augments LLMs with a structured pipeline that bridges natural-language problem descriptions and executable ZK programs. It consists of three main stages: ➀ constraint formulation, ➁ sketch-guided analysis and retrieval, and ➂ interactive generation and repair.
@@
The first stage of ZK-Coder translates natural-language requirements into a structured specification of verification constraints. Drawing on the success of planning-based code generation (Wang et al., 2024; Deng et al., 2024; Li et al., 2023; Jiang et al., 2024b; Wen et al., 2024b), we introduce the ZK Sketch Language (ZKSL), a Python-embedded Domain Specific Language (eDSL) designed to bridge the gap between high-level intent and low-level circuit implementation.
ZKSL is architected to allow LLMs to formalize constraint logic using familiar Pythonic syntax. This isolates the mathematical “what” from the DSL-specific “how” by requiring the explicit modeling of constraints over typed inputs to mirror the verification nature of ZK programs. As formally defined [^1] We also provide concrete running examples in the Supplementary Material. in Fig. 8, a ZKSL Program is structured around a verify function that accepts these typed inputs and executes a sequence of constraint-declaration statements via engine.add($\Phi$). These formulas are composed of relational atoms that correspond to the algebraic primitives identified in § 3.1.2. Crucially, all operations within ZKSL are defined over finite fields (modulo $p$) to faithfully reflect the mathematical foundation of ZK systems. This enables LLMs to express complex verification relations using standardized constructs, serving as a direct logical precursor to the final target code. It allows LLMs to decouple algebraic reasoning from the implementation-level details (e.g., signal wiring and constraint encoding) of specific target DSLs.
ZK-Coder prompts the LLM to generate these constraint specification sketches directly from problem descriptions. To ensure the integrity of the sketch, the framework performs automated syntactic and type/arity consistency checks. If a sketch is found to be invalid, the error trace is fed back to the model for iterative refinement. By serving as a logical representation, ZKSL decouples the complex task of algebraic constraint reasoning from the low-level specifics of ZK DSLs (e.g., signal wiring and constraint specification formats), reducing the overall generation complexity.
4.2.2. Sketch-guided Analysis and Retrieval.
RAG is a powerful technique for grounding LLM generation in low-resource or domain-specific contexts (Joel et al., 2024). ZK-Coder utilizes the ZKSL sketch as a structural query to bridge the gap between abstract constraint specifications and code implementations. Since a ZK program is defined by a set of verification constraints composed of atomic algebraic primitives, the framework traverses the sketch to extract each relation declared via engine.add($\cdot$). To ensure precise retrieval, operators are canonicalized by normalizing aliases (e.g., >= becomes GreaterEqThan) and distinguishing variants based on their arity and field-typing (e.g., differentiating XOr for Booleans vs. Fields). This stage isolates the constituent units within each constraint to determine the precise implementation hints required by the target DSL.
To support this stage, we curate a knowledge base of verified implementations and idiomatic usage patterns for the algebraic primitives defined in § 3.1.2. Instead of text-similarity ranking, ZK-Coder then performs exact-match retrieval for each identified primitive. These patterns are emitted in the original ZKSL specification order to preserve the program’s logical flow during final code generation. This structured retrieval effectively grounds the generation process, collapsing the search space and significantly curbing the model’s tendency to hallucinate incorrect constraint semantics or signal-handling logic.
4.2.3. Interactive Refinement Loop.
While the sketch provides a logical blueprint, our findings in § 3.4 and § 3.5 reveal a persistent gap: many generated programs successfully satisfy DSL grammar rules but fail to faithfully encode the intended verification relation. Unlike mainstream software where errors often manifest as runtime crashes, ZK programs frequently yield under-constrained circuits or over-constrained systems, which silently accept invalid witnesses or sometimes reject valid ones. To bridge this gap, we implement an Interactive Refinement Loop as follows.
  1. (1)
    Constraint Construction Refinement. The candidate program is first submitted to the target compiler (e.g., Circom or Noir) to address DSL-specific failures identified in our study, such as signal mismanagement, illegal signal declarations in control flow, or constraint-encoding violations. We extract these diagnostic traces to iteratively refine the constraint wirings until they’re valid or reach the retry iteration threshold $N_{1}$ ($N_{1}$ is set as 8 in our experiments).
  2. (2)
    Relation Faithfulness Refinement. Once compilable, the framework validates constraint faithfulness. The agent generates representative “Acceptance” and “Rejection” test cases to verify that the circuit correctly enforces the intended relation, while evaluation test cases remain hidden to prevent data-leakage. If any test case fails, the counterexample is fed back to re-examine the algebraic constraints. This forms another refinement loop, and we set the retry threshold $N_{2}$ as 3 in our experiments.
The above processes repeat until the program passes all functional checks or the budgets are exhausted, practically converting failures into guided refinements without manual intervention.
@@ 5. Evaluation
We evaluate ZK-Coder on the task of generating zero-knowledge programs from natural language specifications. Our evaluation is structured around the following research questions, with a focus on end-to-end generation. To clarify, we do not re-evaluate language knowledge (RQ1) or primitive-level coding (RQ2): the former is established in § 3.3, and the latter is handled by our sketching and retrieval mechanism in § 4. Nevertheless, we include an ablation study (RQ5) to quantify the contribution of each component.
@@
  • RQ6 (Robust Generalization): How well does ZK-Coder generalize to novel, complex computational tasks and real-world production-grade ZKP implementations?
  • RQ7 (Failure Analysis): When ZK-Coder fails, what are common failure patterns, and how might they inform future research?
@@ 5.1. Experimental Setup
Our experiments are organized according to the research questions outlined above. For RQ4 (End-to-End Performance), we evaluate on the adapted HumanEval benchmark from ZK-Eval, using the same four models from our measurement study. This comparison establishes how well ZK-Coder outperforms baseline methods in generating correct and provable ZK programs.
For RQ5 (Ablation Study), we conduct an ablation study on all four models, selectively disabling each design component of ZK-Coder, including the sketch language, constraint-guided retrieval, and the interactive repair loop. This aims to quantify their individual impact on overall performance.
@@
All experiments are conducted under a fixed repair budget (see § 4.2.3). The evaluation is conducted on a Linux (Ubuntu 22.04 LTS) server equipped with 256 GiB of RAM.
@@ 5.2. RQ4: End-to-End Performance
We evaluate ZK-Coder’s end-to-end performance. To the best of our knowledge, since no prior work directly addresses ZK program generation from natural language, we establish a simple but meaningful baseline: prompting the model to generate ZK code directly from task descriptions, with a summary of the target DSL’s grammar and few-shot examples. This setting captures the raw capability of LLMs as a lower bound. We provide more baselines in § 5.3.
We meticulously document the performance for the pipeline in three key stages: (1) Sketch correctness measures whether the initial constraint sketch (in ZKSL) produced from the natural language description is correct by passing all hidden test cases; (2) Repair pass rate is conditioned on all correctly generated sketches, and reports the fraction of these that can be successfully repaired by the interactive repair loop; (3) Program correctness is conditioned on all successfully repaired programs, and measures the proportion that pass all hidden test cases. The final overall accuracy aggregates across the full pipeline, reflecting end-to-end program generation accuracy. In addition, we report the average token cost for each model, which quantifies the spending.
@@
Circom GPT-o4-mini 94.12% 91.72% 96.25% 83.38% 2350.31 17.35% 436.98
GPT-o3 97.30% 94.41% 95.64% 87.85% 2155.39 20.29% 433.35
Deepseek V3 72.15% 92.77% 74.10% 49.60% 2958.72 7.35% 325.28
Qwen 3 68.12% 79.07% 78.78% 42.43% 4979.72 4.90% 162.83
Noir GPT-o4-mini 94.12% 97.97% 96.38% 88.87% 648.95 32.21% 152.18
GPT-o3 97.94% 99.85% 100.00% 97.79% 750.63 28.38% 191.72
Deepseek V3 74.88% 93.36% 94.31% 68.66% 628.69 16.18% 100.91
Qwen 3 66.84% 91.88% 89.14% 54.74% 1260.43 11.27% 104.24
Table 2 summarizes the comparative performance of ZK-Coder. Across all tested models and languages, ZK-Coder achieves substantial gains over the baseline. On Circom, GPT-o3 leads with 87.85% accuracy, while open-weight models (DeepSeek V3, Qwen 3) improve from below 10% baseline to 49.60% and 42.43%, respectively. On Noir, success rates are higher overall, likely due to its Rust-like syntax being better aligned with pretraining data. GPT-o3 still leads with 97.79% accuracy, and GPT-o4-mini achieves 88.87%. Open-weight models also show strong gains: Qwen 3 improves from 11.27% to 54.74%, and DeepSeek V3 from 16.18% to 68.66%.
The per-stage breakdown highlights the robustness of the pipeline. OpenAI models show exceptional “zero-shot” capabilities in sketching constraints, while open-weight models exhibit lower correctness rates, suggesting that the leap from natural language to abstract constraint logic remains a challenge for these models. The repair loop maintains strong success rates, exceeding 90% in most settings, though Qwen 3 on Circom achieves 79.07%, indicating some difficulty in recovering from initial errors. Final program correctness is high across most configurations, though DeepSeek V3 and Qwen 3 show comparatively lower rates on Circom, suggesting challenges in implementing correct constraint logic. Overall, the integration of sketching, repair, and semantic validation yields consistently executable and provable programs across the majority of settings.
We also note that ZK-Coder incurs a higher token cost compared to the baseline, which is expected since our agentic pipeline includes sketching, RAG and interactive repair. However, the cost remains modest in practice, with each task requiring fewer than 5,000 tokens on average, corresponding to less than 0.1 USD per task, making the gains in correctness highly cost-effective.
@@ 5.3. RQ5: Ablation Study
Figure 9. Accuracy Comparison of ZK-Coder Ablation Variants (Averaged Across Models).
@@
  • No RAG (V1): Disables sketch-guided RAG for algebraic primitives, forcing the model to rely solely on its internal knowledge.
  • No Sketch (V2): Removes the intermediate ZK constraint sketching and directly prompts the model to output full programs with RAG and the interactive repair loop.
  • No Compile-Repair (V3): Eliminates the interactive repair loop for compiler-identified errors.
  • No Execute-Repair (V4): Retains compiler-guided repair but disables execution and repair.
  • Similarity-based RAG (V5): Replaces our sketch-guided retrieval with a naive text-embedding similarity retrieval on the task description.
  • Only Repair (V6): Omits both retrieval and sketching, relying exclusively on test-and-repair.
Figure 9 presents the Pass@1 accuracy for each ablation variant across Circom and Noir. The results show that every component plays a critical role in the overall success of ZK-Coder. Starting from a full-system average of 65.66% on Circom and 76.87% on Noir, we observe that removing RAG on primitives (V1) causes a significant drop to 53.04% and 72.00% respectively. Disabling the constraint sketching stage (V2) causes further degradation (46.37% on Circom, 64.63% on Noir), highlighting the value of sketching as a scaffold for task understanding and constraint modeling. The most severe drop arises from removing compile-repair (V3), with accuracy falling to 26.04% on Circom and 27.12% on Noir, demonstrating that interactive repair is indispensable. Disabling execute-repair (V4) also leads to measurable reductions (64.48% on Circom, 76.00% on Noir), confirming that semantic validation contributes to robustness. Substituting primitive-in-sketch-guided retrieval with a text-similarity-based strategy (V5) yields weaker performance (51.42% on Circom, 74.59% on Noir), suggesting that sketch-guided retrieval is more effective particularly for Circom. Finally, “Only Repair” (V6) achieves only 43.31% on Circom and 59.80% on Noir, showing that iterative refinement alone cannot overcome the “computation-to-verification” gap. Compared with V6, the full system achieves an up to 22.35% performance leap with modest token cost (20.46%-29.39%), confirming that constraint sketching and template retrieval outperform blind repair cycles.
Answer to RQ5: The ablation study shows that the strong performance of ZK-Coder relies on the synergy of sketching, sketch-guided RAG, and interactive repair. Removing or weakening any of these components causes noticeable accuracy drops.